# Recent Progress in Flavor Model Building

### v1

@@ 3.2 Model Building to Control New Physics Flavor Violation
From a bottom-up point of view, an attractive but fairly restrictive option is the assumption that there are no new sources of flavor (and CP) violation beyond the SM Yukawa couplings. This is known as Minimal Flavor Violation (MFV) [51, 52, 53, 54]. In an MFV scenario, NP contributions are controlled by the same hierarchical structure that suppresses FCNCs in the SM. To implement this idea systematically, the SM Yukawas are promoted to spurion fields with appropriate transformation properties under the flavor symmetry to formally restore flavor invariance of the Yukawa Lagrangian. The SM flavor symmetry group is defined by the fermion kinetic terms
@@
Models based on the $U(2)$ flavor symmetries have been originally proposed in the context of SUSY [21, 22] and later formalized in a general context [56, 57]. The framework of “general MFV” discussed in [26] shares many similar aspects. From a phenomenological point of view, $U(2)^{5}$ models are less restrictive than MFV. While details depend on which spurions break the symmetry, one typically still finds strong correlations between $b\to s$ and $b\to d$ transitions.
@@
While flavor alignment suppresses flavor off-diagonal couplings, one can expect $O(1)$ flavor non-universality. It is important to note that flavor alignment cannot be exact for new physics couplings of left-handed quark doublets unless the couplings are entirely flavor universal. Already in the SM, there is some misalignment between the left-handed up quarks and left-handed down quarks, as parameterized by the CKM matrix. A new non-universal flavor coupling can be aligned either with the left-handed up quarks or with the left-handed down quarks, but not with both of them simultaneously. An irreducible amount of flavor mixing of an order of the Cabibbo angle is predicted either for strange-to-down or charm-to-up transitions. Combining information from neutral kaon mixing and neutral $D$ meson mixing typically gives the strongest constraints on models with flavor alignment [59].
@@ 4.1 Revisiting $U(2)$ and $SU(2)$ Symmetries
Applying singular value decomposition, $Y_{f}=L_{f}\hat{Y}_{f}R_{f}^{\dagger}$, the right-handed rotations $R_{u,d}$ are matrices with $\mathcal{O}(1)$ elements that transform $Y_{u,d}$ into an upper triangular form. In contrast, the left-handed rotations $L_{u,d}$ are perturbative matrices with mixing angles of $\mathcal{O}(b/a)$, $\mathcal{O}(a)$, and $\mathcal{O}(b)$ for the $1$-$2$, $2$-$3$, and $1$-$3$ sectors, respectively. The CKM mixing matrix $V$ is then given by the product of left-handed rotations, $V=L_{u}^{\dagger}L_{d}$. The singular values of both $\hat{Y}_{u}$ and $\hat{Y}_{d}$ are hierarchically ordered as $\{b,a,1\}$. By choosing parameters such that $1\gg a\gg b\gg a^{2}$, this grossly reproduces the observed hierarchies in quark masses between generations as well as the CKM mixing hierarchy. The biggest outlier to this picture is the bottom Yukawa coupling, which has a value of $y_{b}\sim 10^{-2}$.
@@
The IR features described above arise from specific UV dynamics. To illustrate one possible UV completion, consider gauging an $SU(2)$ flavor subgroup, which is anomaly-free in certain cases, such as $q+l$. The SM gauge-singlet flavon field, $\Phi$, is a flavor doublet whose VEV breaks the gauged flavor symmetry at a high-energy scale $\gtrsim$ PeV to comply with the bounds from cLFV [60]. The light Yukawa couplings emerge from a dimension-5 operator in the symmetry-restored EFT, such as $\frac{1}{\Lambda}\bar{q}\Phi Hd_{p}$. When $\langle\Phi\rangle=(0,v_{\Phi})^{T}$, this operator generates the Yukawa coupling for one family, while insertions of $\tilde{\Phi}=\epsilon\Phi^{*}$ provide masses for the other family. With minimal field content completing the operator—for instance, a single copy of a vectorlike quark weak doublet and a flavor singlet—the two operators become proportional, leading to Yukawa matrices with rank two. Consequently, the minimal field content enforces an accidental $U(1)$ flavor symmetry for the first family at tree level, naturally establishing the $1$-$2$ hierarchy. A compelling mechanism to generate masses for the first family is through radiative corrections [60]. Rather than introducing new UV states, new IR states $S$ could run in the loop, with the first-family masses exhibiting only logarithmic sensitivity to their mass scale $m_{S}$, provided $m_{S}\lesssim m_{\Psi}$. This approach fits well with minimal quark-lepton unification, further enhancing the elegance of the mechanism [61].
@@ 4.2 Flavor Deconstruction
This hierarchical structure, with $\langle\phi_{1}\rangle\gg\langle\phi_{2}\rangle>\langle H\rangle$, effectively separates the dynamics of the light and heavy families. The Higgs field is predominantly exposed to the third family dynamics [69, 70, 71, 72] while being effectively shielded from the higher energy scales associated with the lighter families. This framework allows for the TeV-scale NP with built-in protection against FCNCs. Moreover, the multiscale structure preserves the stability of the electroweak scale, ensuring the finite naturalness of the Higgs mass [62]. In essence, this setup elegantly connects the flavor puzzles of both the SM and NP to the electroweak hierarchy problem. The IR phenomenology is governed by the $U(2)^{5}$ flavor symmetry. These models predict large effects in the third generation, specifically $B$-hadron and $\tau$ decays. Direct LHC searches for new resonances coupling mostly to the third generation are less constraining than those for the flavor-universal $U(3)^{5}$ NP due to suppressed production from valance parton distributions and difficulties detecting the third generation in the final state [73, 74, 75, 76]. This is most clearly exemplified by the constraints on semileptonic interactions derived from high-mass Drell-Yan processes [77].
@@ 4.3 Revisiting $U(1)$ Symmetries
As discussed in section 3.1.1, Froggatt-Nielsen models address the flavor problem by assigning flavor-dependent charges to the SM fermions under a $U(1)$ flavor symmetry. This mechanism ensures that masses and mixing angles are parametrically determined by a charge-dependent power of a small spurion, $\epsilon^{n}$. A notable example of such a framework was proposed long ago [16], using $\epsilon\simeq\lambda\simeq 0.2$ to reproduce the observed hierarchical masses and CKM mixing angles successfully. However, it is important to emphasize that the choice of charge assignments is not unique. A wide range of viable configurations exists, as free $O(1)$ parameters can be tuned to accommodate observations with different sets of charges or alternative values of the expansion parameter $\epsilon$. As is often the case, there is a trade-off between the simplicity of charge assignments and the degree of variation allowed in the $O(1)$ parameters. In [81, 82], numerous viable charge assignments have been identified, including some with remarkably minimal charges. Analogously, in the framework of discrete FN models based on the $\mathbb{Z}_{N}$ group, the simplest realizations occur for small values of $N\geq 4$ [83]. These assignments produce distinct deviations in flavor-changing processes, enabling future precision measurements to distinguish between different classes.
@@ 4.5 Flavor Symmetries and Axions
As discussed in section 3.1.1, the SM flavor puzzle can be addressed by a spontaneously broken horizontal $U(1)$ flavor symmetry à la Froggatt-Nielsen. Such a dynamical structure implies observable consequences. In particular, if the $U(1)$ symmetry is not gauged, its spontaneous breaking gives a pseudo-Nambu-Goldstone boson, which is naturally light and can be searched for in the experiment. Furthermore, if the $U(1)$ flavor symmetry is anomalous under QCD, the Goldstone can act as an axion and solve the strong CP problem [100]. In this context, the pseudo-Nambu-Goldstone boson is sometimes referred to as “axiflavon” [101] or “flaxion” [102], see also [103, 104, 105, 106]. Generation-dependent $U(1)$ charges will, in general, give rise to flavor-changing couplings of the axion to fermions. The couplings are proportional to the difference in $U(1)$ charges and suppressed by the axion decay constant $f_{a}$.
@@ 4.7 Modular and Eclectic Flavor Symmetries
Further developments in this direction of flavor model building are eclectic flavor symmetries [142, 143] that combine in a non-trivial way modular flavor symmetries, “traditional” discrete flavor symmetries and CP-like symmetries in top-down string motivated models.
@@ 5 CONCLUSIONS
Flavor models that are based on the popular $U(2)$ or $SU(2)$ flavor symmetries, as well as $U(1)$ models à la FN, continue to evolve and undergo refinement. Flavor models with a spontaneously broken global $U(1)$ symmetry predict the existence of light pseudo-Nambu-Goldstone bosons, such as axions or axion-like particles, characterized by a distinctive pattern of flavor-violating couplings. Developing innovative search strategies for these light, weakly coupled particles remains a highly active area of research.
The concept of flavor deconstruction has recently garnered significant interest. This framework proposes that gauge interactions are non-universal in the UV, with the SM flavor hierarchies emerging from hierarchical scales associated with each family. This setup also addresses the NP flavor puzzle, allowing for the TeV-scale physics, as anticipated by the Higgs hierarchy problem.
Flavor symmetries have been systematically applied to the SMEFT, offering a robust bottom-up framework to explore flavorful new physics. This approach enables a structured investigation of how flavor dynamics can manifest in extensions of the SM. Among recent developments, flavor models based on the clockwork mechanism proposed as an alternative to FN models generate hierarchies through a structured chain of interactions. Additionally, modular flavor symmetries have emerged as a powerful tool in model building, particularly in the neutrino sector, where they offer elegant explanations for the observed flavor patterns by linking flavor symmetries to the geometry of extra dimensions. Finally, innovative concepts such as higher-form symmetries and non-invertible symmetries are beginning to shape the landscape of flavor model building.
@@
The authors are not aware of any affiliations, memberships, funding, or financial holdings that might be perceived as affecting the objectivity of this review.

### v2

@@ 3.2 Model Building to Control New Physics Flavor Violation
From a bottom-up point of view, an attractive but fairly restrictive option is the assumption that there are no new sources of flavor (and CP) violation beyond the SM Yukawa couplings. This is known as the Minimal Flavor Violation (MFV) hypothesis [53, 54, 55, 56]. In an MFV scenario, NP contributions are controlled by the same hierarchical structure that suppresses FCNCs in the SM. To implement this idea systematically, the SM Yukawas are promoted to spurion fields with appropriate transformation properties under the flavor symmetry to formally restore flavor invariance of the Yukawa Lagrangian. The SM flavor symmetry group is defined by the fermion kinetic terms
@@
Models based on the $U(2)$ flavor symmetries have been originally proposed in the context of SUSY [23, 24] and later formalized in a general context [58, 59]. The framework of “general MFV” discussed in [28] shares many similar aspects. From a phenomenological point of view, $U(2)^{5}$ models are less restrictive than MFV models. While details depend on which spurions break the symmetry, one typically still finds strong correlations between $b\to s$ and $b\to d$ transitions.
@@
While flavor alignment suppresses flavor off-diagonal couplings, one can expect $\mathcal{O}(1)$ flavor non-universality. It is important to note that flavor alignment cannot be exact for new physics couplings of left-handed quark doublets unless the couplings are entirely flavor universal. Already in the SM, there is some misalignment between the left-handed up quarks and left-handed down quarks, as parameterized by the CKM matrix. A new non-universal flavor coupling can be aligned either with the left-handed up quarks or with the left-handed down quarks, but not with both of them simultaneously. An irreducible amount of flavor mixing of an order of the Cabibbo angle is predicted either for strange-to-down or charm-to-up transitions. Combining information from neutral kaon mixing and neutral $D$ meson mixing typically gives the strongest constraints on models with flavor alignment [61].
@@ 4.1 Revisiting $U(2)$ and $SU(2)$ Symmetries
Applying singular value decomposition, $Y_{f}=L_{f}\hat{Y}_{f}R_{f}^{\dagger}$, the right-handed rotations $R_{u,d}$ are matrices with $\mathcal{O}(1)$ elements that transform $Y_{u,d}$ into an upper triangular form. In contrast, the left-handed rotations $L_{u,d}$ are perturbative matrices with mixing angles of $\mathcal{O}(b/a)$, $\mathcal{O}(a)$, and $\mathcal{O}(b)$ for the $1$-$2$, $2$-$3$, and $1$-$3$ sectors, respectively. The CKM mixing matrix $V$ is then given by the product of left-handed rotations, $V=L_{u}^{\dagger}L_{d}$. The singular values of both $\hat{Y}_{u}$ and $\hat{Y}_{d}$ are hierarchically ordered as $\{b,a,1\}$. By choosing parameters such that $1\gg a\gg b\gg a^{2}$, this broadly reproduces the observed hierarchies in quark masses between generations as well as the CKM mixing hierarchy. The biggest outlier to this picture is the bottom Yukawa coupling, which has a value of $y_{b}\sim 10^{-2}$.
@@
The IR features described above arise from specific UV dynamics. To illustrate one possible UV completion, consider gauging an $SU(2)$ flavor subgroup, which is anomaly-free in certain cases, such as $q+l$. The SM gauge-singlet flavon field, $\Phi$, is a flavor doublet whose VEV breaks the gauged flavor symmetry at high energies (PeV scale and above) to comply with the bounds from cLFV [62, 63]. The light Yukawa couplings emerge from a dimension-5 operator in the symmetry-restored EFT, such as $\frac{1}{\Lambda}\bar{q}\Phi Hd_{p}$. When $\langle\Phi\rangle=(0,v_{\Phi})^{T}$, this operator generates the Yukawa coupling for one family, while insertions of $\tilde{\Phi}=\epsilon\Phi^{*}$ provide masses for the other family. With minimal field content completing the operator—for instance, a single copy of a vectorlike quark weak doublet and a flavor singlet—the two operators become proportional, leading to Yukawa matrices with rank two. Consequently, the minimal field content enforces an accidental $U(1)$ flavor symmetry for the first family at tree level, naturally establishing the $1$-$2$ hierarchy. A compelling mechanism to generate masses for the first family is through radiative corrections [62]. Rather than introducing new UV states, new IR states $S$ could run in the loop, with the first-family masses exhibiting only logarithmic sensitivity to their mass scale $m_{S}$, provided $m_{S}\lesssim m_{\Psi}$. This approach fits well with minimal quark-lepton unification, further enhancing the elegance of the mechanism [64].
@@ 4.2 Flavor Deconstruction
This hierarchical structure, with $\langle\phi_{1}\rangle\gg\langle\phi_{2}\rangle>\langle H\rangle$, effectively separates the dynamics of the light and heavy families. The Higgs field is predominantly exposed to the third family dynamics [72, 73, 74, 75] while being effectively shielded from the higher energy scales associated with the lighter families. This framework allows for the TeV-scale NP with built-in protection against FCNCs. Moreover, the multiscale structure preserves the stability of the electroweak scale, ensuring the finite naturalness of the Higgs mass [65]. In essence, this setup elegantly connects the flavor puzzles of both the SM and NP to the electroweak hierarchy problem. The IR phenomenology is governed by the $U(2)^{5}$ flavor symmetry. These models predict large effects in the third generation, specifically $B$-hadron and $\tau$ decays. Direct LHC searches for new resonances coupling mostly to the third generation are less constraining than those for the flavor-universal $U(3)^{5}$ NP due to suppressed production from valence parton distributions and difficulties detecting the third generation in the final state [76, 77, 78, 79]. This is most clearly exemplified by the constraints on semileptonic interactions derived from high-mass Drell-Yan processes [80].
@@ 4.3 Revisiting $U(1)$ Symmetries
As discussed in section 3.1.1, Froggatt-Nielsen models address the flavor problem by assigning flavor-dependent charges to the SM fermions under a $U(1)$ flavor symmetry. This mechanism ensures that masses and mixing angles are parametrically determined by a charge-dependent power of a small spurion, $\epsilon^{n}$. A notable example of such a framework was proposed long ago [17], using $\epsilon\simeq\lambda\simeq 0.2$ to reproduce the observed hierarchical masses and CKM mixing angles successfully. However, it is important to emphasize that the choice of charge assignments is not unique. A wide range of viable configurations exists, as free $\mathcal{O}(1)$ parameters can be tuned to accommodate observations with different sets of charges or alternative values of the expansion parameter $\epsilon$. As is often the case, there is a trade-off between the simplicity of charge assignments and the degree of variation allowed in the $\mathcal{O}(1)$ parameters. In [84, 85], numerous viable charge assignments have been identified, including some with remarkably minimal charges. Analogously, in the framework of discrete FN models based on the $\mathbb{Z}_{N}$ group, the simplest realizations occur for small values of $N\geq 4$ [18]. These assignments produce distinct deviations in flavor-changing processes, enabling future precision measurements to distinguish between different classes.
@@ 4.5 Flavor Symmetries and Axions
As discussed in section 3.1.1, the SM flavor puzzle can be addressed by a spontaneously broken horizontal $U(1)$ flavor symmetry à la Froggatt-Nielsen. Such a dynamical structure implies observable consequences. In particular, if the $U(1)$ symmetry is not gauged, its spontaneous breaking gives a pseudo-Nambu-Goldstone boson, which is naturally light and can be searched for in the experiment. Furthermore, if the $U(1)$ flavor symmetry is anomalous under QCD, the Goldstone can act as an axion and solve the strong CP problem [102]. In this context, the pseudo-Nambu–Goldstone boson is sometimes referred to as an “axiflavon” [103] or “flaxion” [104], see also [105, 106, 107, 108]. Generation-dependent $U(1)$ charges will, in general, give rise to flavor-changing couplings of the axion to fermions. The couplings are proportional to the difference in $U(1)$ charges and suppressed by the axion decay constant $f_{a}$.
@@ 4.7 Modular and Eclectic Flavor Symmetries
Further developments in this direction of flavor model building are eclectic flavor symmetries [144, 145] that combine in a nontrivial way modular flavor symmetries, “traditional” discrete flavor symmetries, and CP-like symmetries in top-down string motivated models.
@@ 5 CONCLUSIONS
Flavor models that are based on the popular $U(2)$ or $SU(2)$ flavor symmetries, as well as $U(1)$ models à la Froggatt-Nielsen (FN), continue to evolve and undergo refinement. Flavor models with a spontaneously broken global $U(1)$ symmetry predict the existence of light pseudo-Nambu-Goldstone bosons, such as axions or axion-like particles, characterized by a distinctive pattern of flavor-violating couplings. Developing innovative search strategies for these light, weakly coupled particles remains a highly active area of research.
The concept of flavor deconstruction has recently garnered significant interest. This framework proposes that gauge interactions are non-universal in the UV, with the SM flavor hierarchies emerging from hierarchical scales associated with each family. This setup also addresses the NP flavor puzzle, allowing for physics at the TeV scale, as anticipated by the Higgs hierarchy problem.
Flavor symmetries have been systematically applied to the SMEFT, offering a robust bottom-up framework to explore flavorful new physics. This approach enables a structured investigation of how flavor dynamics can manifest in extensions of the SM. Among recent developments, flavor models based on the clockwork mechanism proposed as an alternative to FN models generate hierarchies through a structured chain of interactions. Additionally, modular flavor symmetries have emerged as a novel tool in model building, particularly in the neutrino sector, where they offer elegant explanations for the observed flavor patterns by linking flavor symmetries to the geometry of extra dimensions. Finally, innovative concepts such as higher-form symmetries and non-invertible symmetries are beginning to shape the landscape of flavor model building.
@@
The authors a.re not aware of any affiliations, memberships, funding, or financial holdings that might be perceived as affecting the objectivity of this review.